# Kendall’s $\tau$ Coefficient for Logits Distillation

### v1

@@ 6 Conclusion
In this paper, we investigate the optimization process of logit distillation and identify that the Kullback-Leibler divergence tends to overlook the knowledge embedded in the smaller channels of the output. Moreover, KL divergence does not guarantee alignment between the classification results of the student and teacher models. To address this issue, we introduce a ranking loss that encourages the student model to pay more attention to the knowledge contained within the smaller channels, while also enforcing alignment with the teacher’s predictive outcomes. We provide a theoretical analysis demonstrating that ranking loss can serve as an effective auxiliary loss to KL loss. Building on the Kendall $\tau$ coefficient, we propose three differential loss forms tailored for the distillation context. Extensive experiments validate that our approach significantly enhances the distillation performance across various teacher-student architectures. As a plug-and-play auxiliary method, it also effectively improves the efficacy of logit-based distillation methods.
Appendix A Appendix / supplemental material
A.1 Limitations
We have not explored the effectiveness of our method on other downstream tasks.
A.2 KD for Transformer
We use the AdamW optimizer and train for 300 epochs with an initial learning rate of 5e-4 and a weight decay of 0.05. The minimum learning rate is 5e-6, and the patch size is 16. We set $\alpha$ = 1, $\beta$ = 1, $\gamma$ = 0.5, and batch size is 128. The GPU we used is a single RTX4090.
A.3 Ablation of Temperature
There has been extensive and detailed research on the temperature of the KL loss (e.g., CTKD[20], MLKD[14], LSKD[27]), achieving significant results. Unlike these studies, our approach aims to move away from focusing on the KL divergence and instead explore plug-and-play auxiliary losses to guide the KL divergence. Therefore, in our experiments, the temperature and other parameters of the KL divergence are kept the same as those in the respective baselines (with dynamic/multi-level temperatures used in CTKD+Ours and MLKD+Ours). The proposed ranking loss does not have a temperature parameter because the temperature does not affect the order of logits. Instead, the control over the distribution can be achieved by manipulating the steepness parameter $k$ of the sign function, and ablation experiments for $k$ are shown in Tab. 8. Meanwhile, although the temperature helps control the logit distribution, additional measures are needed to prevent the KL divergence from falling into suboptimal. We conducted an ablation study on the temperature, as shown in Tab. 9. The results demonstrate that ranking loss as a plug and play loss can bring further improvements under multiple temperature settings.
Tempetature T = 4 T = 5 T = 6 T = 10
KD 73.33 73.39 73.43 73.55
KD+RKKD 73.56(+0.23) 73.49(+0.10) 74.36(+0.0.93) 74.04(+0.49)
A.4 Ablation of Normalization
@@
Table 10: Ablation of Normalization+.
@@
KD+RKKD w/o Norm 76.38(+2.31) 74.07(+0.53)
KD+RKKD w Norm 75.98(+1.91) 74.49(+0.95)

### v2

@@ 6 Conclusion
In this paper, we investigate the optimization process of logit distillation and identify that the Kullback-Leibler divergence tends to overlook the knowledge embedded in the low-probability channels of the output. Moreover, KL divergence does not guarantee alignment between the classification results of the student and teacher models. To address this issue, we introduce a plug-and-play ranking loss based on Kendall’s $\tau$ Coefficient that encourages the student model to pay more attention to the knowledge contained within the low-probability channels, while also enforcing alignment with the teacher’s predictive outcomes. We provide a theoretical analysis demonstrating that the gradients of the ranking loss are less affected by channel scale and that its optimization objective is consistent with that of KL divergence, making it an effective auxiliary loss for distillation. Extensive experiments validate that our approach significantly enhances the distillation performance across various datasets and teacher-student architectures.
Acknowledgements
This work is supported by the National Key R&D Program of China (2022YFB4701400/4701402), SSTIC Grant (KJZD20230923115106012, KJZD20230923114916032, GJHZ20240218113604008), Beijing Key Lab of Networked Multimedia.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Implementation Details for Transformer
We use the AdamW optimizer and train for 300 epochs with an initial learning rate of 5e-4 and a weight decay of 0.05. The minimum learning rate is 5e-6, and the patch size is 16. We set $\alpha$ = 1, $\beta$ = 1, $\gamma$ = 0.5, and batch size is 128. The value range of ranking loss is $[-1,1]$ We use a single RTX4090 for CIFAR-100 and 4 RTX4090 for ImageNet.
A.2 Ablation of Normalization
@@
Table 10: Ablation of Normalization.
@@
KD+Ours w/o Norm 76.38(+2.31) 74.07(+0.53)
KD+Ours w Norm 75.98(+1.91) 74.49(+0.95)
A.3 Different Forms of Ranking Loss
In the initial application of the ranking loss(Zheng et al., 2023), it is necessary to compute the gradients for two input vectors separately. However, In the scenario of distillation, the soft labels from the teacher model do not require gradients. Therefore, we propose three variants of the diff-Kendall ranking loss suitable for distillation scenarios, aimed at further exploring the role of ranking loss in distillation. Since the sample pair $(z^{t}_{i}-z^{t}_{j})$ itself has a sign,, we can derive an equivalent form from Eq. 5:
$\tau=\frac{\sum_{i}\sum_{j<i}sgn((z^{t}_{i}-z^{t}_{j})\cdot(z^{s}_{i}-z^{s}_{j}))}{\frac{1}{2}C(C-1)}$ (13)
For Eq. 5 and Eq. 13, we can similarly transform them into differential forms of ranking loss by replacing $sgn$ with $tanh$, and gradients are not computed for the output part corresponding to the teacher, which is:
$L_{\tau}^{form1}=\frac{\sum_{i}\sum_{j<i}tanh(z^{t}_{i}-z^{t}_{j})_{detach}\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (14)
$L_{\tau}^{form2}=\frac{\sum_{i}\sum_{j<i}tanh[(z^{t}_{i}-z^{t}_{j})_{detach}\cdot(z^{s}_{i}-z^{s}_{j})]}{\frac{1}{2}C(C-1)}$ (15)
Where $()_{detach}$ means not participating in training. Further, since the gradient of the teacher’s output is not required, we can directly use the sign function instead of the $tanh$ function to obtain the ranking loss:
$L_{\tau}^{form3}=\frac{\sum_{i}\sum_{j<i}sgn(z^{t}_{i}-z^{t}_{j})_{detach}\cdot tanh(z^{s}_{i}-z^{s}_{j})}{\frac{1}{2}C(C-1)}$ (16)
Different Forms of Ranking Loss. In our investigation, we examined the performance of the three forms of distillation presented in Sec. A.3, as depicted in Tab. 11. Form1 exhibited the most superior performance, followed by Form2, with Form3 trailing yet still enhancing the efficacy of the original knowledge distillation. This suggests that within the ranking loss, it is imperative to optimize considering the magnitude of the teacher’s logits differences as a coefficient for the loss, rather than merely optimizing as a sign function.
Loss Form KD KD+Form1 KD+Form2 KD+Form3
Similar Structure 72.01 73.47(+1.46) 73.42(+1.41) 73.38(+1.37)
Different Structure 72.74 73.50(+0.76) 73.36(+0.62) 73.05(+0.31)
A.4 Algorithm
Algorithm 1 Plug-and-Play Ranking Loss for Logit Distillation
Input: Transfer set $\mathcal{D}$ with samples of image-label pair $\{x_{n},y_{n}\}_{n=1}^{N}$, base temperature $T$, teacher $f_{t}$, student $f_{s}$, knowledge distillation Loss $\mathcal{L}_{\text{KD}}$, ranking Loss $\mathcal{L}_{\text{RK}}$, the weight of ranking Loss $\gamma$.
Output: Trained student model $f_{s}$
for  $(x_{n},y_{n})$ in $\mathcal{D}$ do
       Get the logits of Teacher and student:$z^{t}=f_{t}(x_{n})$, $z^{s}=f_{s}(x_{n})$
Calculate the probability with temperature: $q^{t}=softmax(\frac{z^{t}}{T})$ , $q^{s}=softmax(\frac{z^{s}}{T})$
Get the normalized logits of Teacher and student: $\hat{z}^{t}=\frac{z^{t}-\bar{z^{t}}}{std(z^{t})}$, $\hat{z}^{s}=\frac{z^{s}-\bar{z^{s}}}{std(z^{s})}$ Update $f_{s}$ towards minimizing: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{KD}}(q^{t},q^{s})+\gamma\cdot\mathcal{L}_{\text{RK}}(\hat{z}^{t},\hat{z}^{s})$
end for
A.5 Derivation of the KL Loss and Ranking Loss
Derivation of the KL Divergence with Respect to Student Logits in Knowledge Distillation.
Denote the teacher’s logits as $\mathbf{z}^{t}=[z^{t}_{1},z^{t}_{2},\dotsc,z^{t}_{C}]$.
Denote the student’s logits as $\mathbf{z}^{s}=[z^{s}_{1},z^{s}_{2},\dotsc,z^{s}_{C}]$.
Let $T$ be the temperature scaling factor used in the softmax function.
Teacher probabilities can be calculated as:
$q^{t}_{i}=\frac{\exp\left(\dfrac{z^{t}_{i}}{T}\right)}{\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{t}_{j}}{T}\right)},\quad\text{for }i=1,2,\dotsc,C.$ (17)
Student probabilities can be calculated as:
$q^{s}_{i}=\frac{\exp\left(\dfrac{z^{s}_{i}}{T}\right)}{\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)},\quad\text{for }i=1,2,\dotsc,C.$ (18)
The loss function used in knowledge distillation is the scaled Kullback-Leibler (KL) divergence between the teacher and student probability distributions:
$L_{\text{KD}}=T^{2}\cdot\mathrm{KL}\left(q^{t}\,\|\,q^{s}\right)=T^{2}\sum\limits_{i=1}^{C}q^{t}_{i}\log\left(\frac{q^{t}_{i}}{q^{s}_{i}}\right).$ (19)
The derivative of the loss with respect to the student logits $z^{s}_{i}$ is:
$\frac{\partial L_{\text{KD}}}{\partial z^{s}_{i}}=-T^{2}\sum\limits_{k=1}^{C}q^{t}_{k}\frac{\partial\log q^{s}_{k}}{\partial z^{s}_{i}}.$ (20)
Taking the natural logarithm:
$\log q^{s}_{k}=\frac{z^{s}_{k}}{T}-\log\left(\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)\right).$ (21)
Compute the partial derivative:
$\frac{\partial\log q^{s}_{k}}{\partial z^{s}_{i}}=\frac{\partial}{\partial z^{s}_{i}}\left(\frac{z^{s}_{k}}{T}\right)-\frac{\partial}{\partial z^{s}_{i}}\left(\log\left(\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)\right)\right).$ (22)
Compute each term separately:
$\frac{\partial}{\partial z^{s}_{i}}\left(\frac{z^{s}_{k}}{T}\right)=\frac{1}{T}\delta_{ik},\quad\delta_{ik}=\begin{cases}1,&\text{if }i=k,\\ 0,&\text{if }i\neq k.\end{cases}\textbf{}$ (23)
$\frac{\partial}{\partial z^{s}_{i}}\left(\log\left(\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)\right)\right)=\frac{1}{\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)}\cdot\frac{\partial}{\partial z^{s}_{i}}\left(\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)\right).$ (24)
The derivative inside the sum is:
$\frac{\partial}{\partial z^{s}_{i}}\left(\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)\right)=\exp\left(\dfrac{z^{s}_{i}}{T}\right)\cdot\frac{1}{T}=\frac{\exp\left(\dfrac{z^{s}_{i}}{T}\right)}{T}.$ (25)
Therefore, the second term becomes:
$\frac{1}{\sum\limits_{j=1}^{C}\exp\left(\dfrac{z^{s}_{j}}{T}\right)}\cdot\frac{\exp\left(\dfrac{z^{s}_{i}}{T}\right)}{T}=\frac{1}{T}q^{s}_{i}.$ (26)
Thus, the total derivative is:
$\frac{\partial\log q^{s}_{k}}{\partial z^{s}_{i}}=\frac{1}{T}\delta_{ik}-\frac{1}{T}q^{s}_{i}=\frac{1}{T}(\delta_{ik}-q^{s}_{i}).$ (27)
We now substitute $\dfrac{\partial\log q^{s}_{k}}{\partial z^{s}_{i}}$ back into the expression for the derivative of the loss:
$\frac{\partial L_{\text{KD}}}{\partial z^{s}_{i}}$ $=-T^{2}\sum\limits_{k=1}^{C}q^{t}_{k}\left(\frac{1}{T}(\delta_{ik}-q^{s}_{i})\right)$ (28)
$=-T\sum\limits_{k=1}^{C}q^{t}_{k}(\delta_{ik}-q^{s}_{i}).$
Considering:
$\sum\limits_{k=1}^{C}q^{t}_{k}\delta_{ik}=q^{t}_{i},$ (29)
$\sum\limits_{k=1}^{C}q^{t}_{k}q^{s}_{i}=q^{s}_{i}\sum\limits_{k=1}^{C}q^{t}_{k}=q^{s}_{i}\cdot 1=q^{s}_{i},\quad\sum\limits_{k=1}^{C}q^{t}_{k}=1.$ (30)
Therefore, the loss derivative simplifies to:
$\frac{\partial L_{\text{KD}}}{\partial z^{s}_{i}}=-T\left(q^{t}_{i}-q^{s}_{i}\right).$ (31)
Derivation of the Ranking Loss with Respect to Student Logits in Knowledge Distillation.
The Rank loss is calculated as:
$L_{RK}=-\frac{\sum_{i}\sum_{j<i}tanh(k(z^{t}_{i}-z^{t}_{j}))\cdot tanh(k(z^{s}_{i}-z^{s}_{j}))}{\frac{C(C-1)}{2}}$ (32)
The derivation is calculated as:
$\frac{\partial L_{RK}}{\partial z^{s}_{i}}$ $=\frac{1}{\frac{C(C-1)}{2}}\sum_{j\neq i}\frac{\partial}{\partial z^{s}_{i}}\frac{1}{2}\cdot\left[\tanh\left(k(z^{s}_{i}-z^{s}_{j})\right)\tanh\left(k(z^{t}_{i}-z^{t}_{j})\right)\right]$ (33)
$=\frac{1}{C(C-1)}\sum_{j\neq i}\frac{\partial}{\partial z^{s}_{i}}\left[\tanh\left(k(z^{s}_{i}-z^{s}_{j})\right)\tanh\left(k(z^{t}_{i}-z^{t}_{j})\right)\right]$ (34)
Denote $\phi_{ij}$ as:
$\phi_{ij}=\tanh\left(k(z^{s}_{i}-z^{s}_{j})\right)\tanh\left(k(z^{t}_{i}-z^{t}_{j})\right)$ (35)
Its derivative w.r.t. $z^{s}_{i}$ is:
$\frac{\partial\phi_{ij}}{\partial z^{s}_{i}}=\left[1-\tanh^{2}\left(k(z^{s}_{i}-z^{s}_{j})\right)\right]\cdot k\cdot\tanh\left(k(z^{t}_{i}-z^{t}_{j})\right)$ (36)
Finally, the gradient of $L_{RK}$ with respect to $z^{s}_{i}$ is:
$\frac{\partial L_{RK}}{\partial z^{s}_{i}}=-\frac{k}{C(C-1)}\sum_{j\neq i}\left[1-\tanh^{2}\left(k(z^{s}_{i}-z^{s}_{j})\right)\right]\tanh\left(k(z^{t}_{i}-z^{t}_{j})\right)$ (37)
A.6 Derivation of the Differential Ranking Loss
In this section, we explain how we get the final form of Eq. 6. Noticed that:
$tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}=\frac{e^{2x}-1}{e^{2x}+1}=1-\frac{2}{e^{2x}+1}$ (38)
Then, we can get the final form of Eq. 6 by expanding the $tanh$ function.
$\tau_{d}$ $=\frac{\sum_{i}\sum_{j<i}tanh(k\cdot(z^{t}_{i}-z^{t}_{j}))\cdot tanh(k\cdot(z^{s}_{i}-z^{s}_{j}))}{\frac{1}{2}C(C-1)}$ (39)
$=\frac{2}{C(C-1)}\cdot\sum_{i}\sum_{j<i}\left(1-\frac{2}{1+e^{2\cdot(z^{t}_{i}-z^{t}_{j})\cdot k}}\right)\cdot\left(1-\frac{2}{1+e^{2\cdot(z^{s}_{i}-z^{s}_{j})\cdot k}}\right)$ (40)
A.7 Does Temperature Solve the Problem of Ignoring Low-Probability Channels
Some methods (Wei et al., 2025; Cheng et al., 2025) address the issue of neglecting low-probability channels by employing temperature scaling. We examined the probability distribution of LSKD (Sun et al., 2024), an approach using adaptive temperature that has shown promising results. As shown in Fig .9, while LSKD somewhat alleviates the problem of low-probability channels, the transformed probabilities still predominantly occupy these low-probability channels. Thus, using temperature scaling does not effectively resolve the issue of neglecting low-probability channels.
Figure 9: probability produced by KD and LSKD
A.8 Combining with Other Correlation Coefficients.
DIST (Huang et al., 2022) explores the use of the Pearson correlation coefficient for knowledge distillation. However, unlike our proposed method, the Pearson correlation coefficient measures the linear relationship between two variables. Our method focuses on the ranking relationship of logits, which is not strictly a linear relationship. Therefore, Kendall’s $\tau$ correlation coefficient may be more suitable for capturing this non-linear ranking information. To further investigate the combined effect of using both linear and non-linear auxiliary functions, we incorporate ranking loss into DIST, which improves its performance. The results are shown in Tab. 12.
Table 12: Experiments on Combining DIST.
Teacher ResNet32$\times$4 ResNet32$\times$4 WRN-40-2 VGG13 WRN-40-2
Student WRN-16-2 WRN-40-2 SHN-V1 VGG8 WRN-40-1
DIST (Huang et al., 2022) 75.58 78.02 76.00 73.80 74.73
DIST+Ours 75.85(+0.27) 78.54(+0.52) 76.23(+0.23) 74.06(+0.26) 74.86(+0.13)
A.9 More ImageNet Results
We further augment CTKD (Li et al., 2023) and NKD (Yang et al., 2023a) with the proposed ranking loss and evaluate them on the ImageNet dataset; the results are reported in Tab. 13.
Table 13: The top-1 and top-5 accuracy (%) on the ImageNet validation set. The teacher and student are ResNet34 and ResNet18
Method Top-1 Acc. Top-5 Acc.
AT (Zagoruyko & Komodakis, 2017) 70.69 90.01
OFD (Heo et al., 2019) 70.81 89.98
CRD (Tian et al., 2019) 71.17 90.13
CTKD (Li et al., 2023) 71.32 90.27
CTKD+Ours 71.68(+0.36) 90.30(+0.03)
NKD (Yang et al., 2023a) 71.96 90.27
NKD+Ours 72.06(+0.10) 90.49(+0.22)